### v1

# Latent Denoising Makes Good Visual Tokenizers

Abstract
Despite their fundamental role, it remains unclear what properties could make visual tokenizers more effective for generative modeling. We observe that modern generative models share a conceptually similar training objective—reconstructing clean signals from corrupted inputs such as Gaussian noise or masking—a process we term denoising. Motivated by this insight, we propose aligning tokenizer embeddings directly with the downstream denoising objective, encouraging latent embeddings to be more easily reconstructed even when heavily corrupted. To achieve this, we introduce the Latent Denoising Tokenizer ($l$-DeTok), a simple yet effective tokenizer trained to reconstruct clean images from latent embeddings corrupted by interpolative noise and random masking. Extensive experiments on ImageNet $256\times 256$ demonstrate that our tokenizer consistently outperforms standard tokenizers across six representative generative models. Our findings highlight denoising as a fundamental design principle for tokenizer development, and we hope it could motivate new perspectives for future tokenizer design. Code is available at: https://github.com/Jiawei-Yang/DeTok
1 Introduction
Modern visual generative models commonly operate on compact latent embeddings, produced by tokenizers, to circumvent the prohibitive complexity of pixel-level modeling [51, 46, 5, 36]. Current tokenizers are typically trained as standard variational autoencoders [33], primarily optimizing for pixel-level reconstruction. Despite their critical influence on downstream generative quality, it remains unclear what properties enable more effective tokenizers for generation. As a result, tokenizer development has lagged behind recent rapid advances in generative model architectures.
In this work, we ask: What properties can make visual tokenizers more effective for generative modeling? We observe that modern generative models, despite methodological differences, share a conceptually similar training objective—reconstructing original signals from corrupted ones. For instance, diffusion models remove diffusion-induced noise to recover clean signals [31, 46], while autoregressive models reconstruct complete sequences from partially observed contexts [36, 5], analogous to removing “masking noise” [30, 11, 15]. We collectively refer to these reconstruction-from-deconstruction processes as denoising.
This unified denoising perspective of modern generative models suggests that effective visual tokenizers for these models should produce latent embeddings that are reconstructable even under significant corruption. Such embeddings naturally align with the denoising objectives of downstream generative models, facilitating their training and subsequently enhancing their generation quality.
Motivated by this insight, we propose to train tokenizers as latent denoising autoencoders, termed as $l$-DeTok. Specifically, we deconstruct latent embeddings via interpolative noise, obtaining corrupted latents by interpolating original embeddings with Gaussian noise. The tokenizer decoder is then trained to reconstruct clean images from these heavily noised latent embeddings. Additionally, we explore random masking, similar to masked autoencoders (MAE) [30], as an alternative form of deconstruction and find it similarly effective.
Conceptually, these deconstruction-reconstruction strategies encourage latent embeddings to be robust, stable, and easily reconstructable under strong corruption, aligning with the downstream denoising tasks central to generative models. Indeed, our experiments show that stronger noise (i.e., strong latent noise and high masking ratio) used in our $l$-DeTok training usually leads to better downstream generative performance.
We demonstrate the effectiveness and generalizability of $l$-DeTok across six representative generative models, including non-autoregressive (DiT [46], SiT [42], LightningDiT [64]) and autoregressive models (MAR [36], RasterAR, RandomAR [36, 44, 68]) on the ImageNet generation benchmark. We find that compared to recent semantics-distilled tokenizers [64, 6], which achieve strong performance on diffusion models but generalize poorly to autoregressive ones, our tokenizer demonstrates significantly broader generalizability. For example, by adopting our tokenizer—without modifying model architectures—we push the limits for MAR models [36], improving FID from 2.31 to 1.55 for MAR-B, matching the performance of the original huge-sized MAR (1.55). For MAR-L, FID improves from 1.78 to 1.35. Importantly, these gains come without semantics distillation, thus avoiding dependencies on visual encoders pretrained at a far larger scale [43, 49].
In summary, our work demonstrates a simple yet crucial insight: explicitly incorporating denoising objectives into tokenizer training significantly enhances their effectiveness for generative modeling since it is downstream task-aligned. We hope this perspective will stimulate new research directions in tokenizer design and accelerate future advances in generative modeling.
2 Related Work
Representation learning in visual recognition.
Representation learning has been a decades-long pursuit in visual recognition, aiming to discover transferable embeddings that generalize across diverse downstream tasks [2]. Starting from supervised representations [17, 24, 65], self-supervised methods have significantly pushed the boundaries of transferability without explicit human annotations [29, 9, 27, 30, 43, 11, 4]. At the core of these approaches lies a foundational principle: pre-training should encourage representations to encode the information most relevant to downstream tasks. This principle has inspired the design of diverse and effective pretext tasks, such as instance discrimination [29, 9, 3], self-distillation [27, 10, 4, 43], and masked-image reconstruction [30, 43, 72], each aligning representations with downstream utility. These insights motivate us to explore tokenizer embeddings that align with downstream generative tasks.
Visual tokenizers for generative modeling.
Modern generative models typically rely on tokenizers to encode images into compact latent embeddings, significantly reducing computational complexity compared to pixel-level modeling [19, 51, 46, 34, 55, 69, 37, 7, 1, 12, 48]. While conventional tokenizers optimize pixel reconstruction with KL-regularization [35], recent approaches [6, 64, 7, 37] have advocated semantics distillation from powerful pretrained vision models [43, 49]. However, such semantics-distilled tokenizers rely inherently on a two-stage pipeline: first training a vision encoder at a significantly larger scale in terms of compute and data, then distilling its features into latent embeddings. Moreover, our experiments reveal a surprising limitation (Table 1): semantics-distilled continuous tokens enhance non-autoregressive diffusion model generation but notably degrade autoregressive methods, questioning the common assumption that tokenizer improvements in one generative paradigm transfer to others. In contrast, our $l$-DeTok generalizes better, consistently benefiting diverse models—yet without relying on semantics distillation.
Generative modeling frameworks.
Generative modeling encompasses various downstream generation frameworks and can be broadly categorized into autoregressive (AR) [8, 19, 5, 36, 59, 21, 44, 68] and diffusion-based non-autoregressive (non-AR) [31, 51, 46, 42, 64] methods. AR models factorize latent representations sequentially, predicting tokens step-by-step conditioned on partially generated contexts, while non-AR models jointly predict all latent tokens through iterative refinement, typically leveraging diffusion [31, 56] or flow-based processes [39, 20]. Despite their methodological diversity, these generative paradigms depend on the quality of latent embeddings produced by upstream tokenizers [28]. This inherent dependence highlights the critical importance of studying latent embeddings: improvements at the tokenizer level have the potential to yield better generative performance across the full spectrum of downstream generative frameworks.
3 Method
Our goal is to design visual tokenizers that are more effective for generative modeling compared to standard autoencoders trained for pixel reconstruction. This section first revisits the core training objective shared by all modern generative models, i.e., denoising, to motivate our design, and then details our latent denoising tokenizers ($l$-DeTok).
3.1 Preliminaries of Generative Modeling
Modern generative frameworks can be primarily divided into non-autoregressive (non-AR) and autoregressive (AR) paradigms. Despite their methodological differences, both paradigms aim to gradually reconstruct the original representations from deconstructed ones.
Non-autoregressive generative models.
Non-autoregressive models, exemplified by diffusion [31, 46] and flow-matching methods [39, 42], learn to iteratively refine latent representations deconstructed by controlled noise. Given a latent representation of an image $\mathbf{X}_{0}$, the forward noising process progressively corrupts these latents into $\mathbf{X}_{t}$:
$\mathbf{X}_{t}=a(t)\,\mathbf{X}_{0}+b(t)\,\boldsymbol{\epsilon}_{t},\quad\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (1)
where $a(t)$ and $b(t)$ are noise schedules. Generative models are trained to revert this deconstruction:
$\mathcal{L}(\boldsymbol{\theta})={{\mathbb{E}_{\mathbf{X},\boldsymbol{\epsilon},t}}{\left[{\left\|{{\epsilon_{\boldsymbol{\theta}}}{\left(\mathbf{X}_{t},t\right)}}-\boldsymbol{\epsilon}_{t}\right\|}^{2}\right]}},$ (2)
where ${\epsilon_{\boldsymbol{\theta}}}$ is a learnable noise estimator parameterized by $\boldsymbol{\theta}$. Essentially, non-AR diffusion models learn to reconstruct original latents from intermediate latents deconstructed by noise.
Autoregressive generative models.
Autoregressive approaches factorize image generation into a sequential prediction problem. Given an ordered sequence of latent tokens $\{\mathbf{x}^{1},\dots,\mathbf{x}^{N}\}$, AR methods factorize the joint distribution as:
$p_{\boldsymbol{\theta}}(\mathbf{x})=\prod_{i=1}^{N}p_{\boldsymbol{\theta}}(\mathbf{x}^{i}|\mathbf{x}^{1},\dots,\mathbf{x}^{i-1}),$ (3)
where $\mathbf{x}^{i}$ denotes the latent tokens generated at step $i$. Recent generalized AR variants extend this framework to arbitrary generation orders [36, 68, 44] or set-wise prediction strategies [59, 50]. Nonetheless, the fundamental training objective remains consistent: reconstructing full sequences from partially observed—or equivalently, partially masked—contexts. In other words, AR models learn to reconstruct original latents from intermediate latents deconstructed by masking.
3.2 Latent Denoising Tokenizers
Figure 1: Our latent denoising tokenizers ($l$-DeTok) framework. During tokenizer training, we randomly mask input patches (masking noise) and interpolate encoder-produced latent embeddings with Gaussian noise (interpolative latent noise). The decoder processes these deconstructed latents and mask tokens to reconstruct the original images in pixels. We refer to this process as denoising. When serving as a tokenizer for downstream generative models, both noises are disabled.
Motivated by the discussions above, we propose latent denoising tokenizer ($l$-DeTok), a simple tokenizer trained by reconstructing original images from deconstructed latent representations. This deconstruction-reconstruction design aligns with the denoising tasks employed by modern generative models. Figure 1 shows an overview of our method. We detail each component next.
Overview.
Our tokenizer follows an encoder-decoder architecture based on Vision Transformers (ViT) [18, 61]. Input images are divided into non-overlapping patches, linearly projected into embedding vectors, and added with positional embeddings. During training, we deconstruct these embeddings using two complementary strategies: (i) injecting noise in latent embeddings and (ii) randomly masking image patches. The decoder reconstructs original images from these deconstructed embeddings. This strategy encourages easy-to-reconstruct latent embeddings under heavy corruption, aiming to simplify downstream denoising tasks in generative models.
Noising as deconstruction.
Our core idea is to deconstruct latent embeddings by noise interpolation. Specifically, given latent embeddings $\mathbf{x}$ from the encoder, we interpolate them with Gaussian noise as follows:
$\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\boldsymbol{\varepsilon}(\gamma),\quad\text{where}\quad\boldsymbol{\varepsilon}(\gamma)\sim\gamma\cdot\mathcal{N}(\mathbf{0},\mathbf{I}),\quad\tau\sim\mathcal{U}(0,1).$ (4)
Here, the scalar $\gamma$ controls noise standard deviation, and the factor $\tau$ specifies noise level. Critically, this interpolative strategy differs from the conventional additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\varepsilon$, employed by standard VAEs [47] or DAEs [62], as it ensures latents can be effectively and heavily corrupted when the noise level $\tau$ is high. Moreover, random sampling of $\tau$ encourages latents to remain robust across diverse corruption levels. Unlike traditional DAEs [62], which apply additive noise directly in pixel space, our latent-space interpolation directly aligns with downstream generative models that operate in latent embedding spaces. At inference time, latent noising is disabled ($\tau=0$).
Masking as deconstruction.
We further generalize our denoising perspective by interpreting masking as another form of latent deconstruction [15, 11, 30]. Inspired by masked autoencoders (MAE) [30], we randomly mask a subset of image patches. Different from MAE, we use a random masking ratio. Concretely, given an input image partitioned into patches, we mask a random subset, where the masking ratio $m$ is sampled from a slightly biased uniform distribution:
$m=\max(0,\mathcal{U}(-0.1,M)),$ (5)
where $\mathcal{U}(-0.1,M)$ denotes a uniform distribution on $[-0.1,M]$. The slight bias towards zero reduces the distribution gap between training and inference (no masking). The encoder processes only the visible patches, and masked positions are represented by shared learnable [MASK] tokens at the decoder input. At inference time, all patches are visible ($m=0$).
Training objectives.
Our decoder reconstructs the original images from corrupted latent embeddings. The training objective follows established practice [51, 19, 69], combining pixel-wise mean-squared-error (MSE), latent-space KL-regularization [47], perceptual losses (VGG- [54] and ConvNeXt-based [40] as in [69]), and an adversarial GAN objective [25]:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}+\lambda_{\text{percep}}\mathcal{L}_{\text{percep}}+\lambda_{\text{GAN}}\mathcal{L}_{\text{GAN}},$ (6)
where each $\lambda$ controls the contribution of the corresponding loss component.
4 Implementation
We describe the implementation details, including datasets, evaluation metrics, tokenizer and generative model training procedures.
Dataset and metrics.
We perform all experiments on ImageNet [52] at $256\times 256$ resolution. Images are center-cropped, randomly horizontally flipped during training, and normalized to $[-1,1]$. We evaluate generative models using Fréchet Inception Distance (FID), Inception Score (IS), precision and recall, following [16]. For tokenizer reconstruction, we report reconstruction-FID (rFID).
Tokenizer baselines.
We benchmark our tokenizer against a diverse set of publicly available tokenizers: (i) MAR-VAE from MAR [36], trained on ImageNet using the implementation from [19]; (ii) VA-VAE [64], aligning latent embeddings with DINOv2 features [43]; (iii) MAETok [6], distilling HOG [13], DINOv2 [43], and CLIP [49] features through auxiliary decoders; (iv) SD-VAE from Stable-Diffusion [51], trained on significantly larger datasets. Additionally, for controlled comparisons, we also train our own baseline tokenizer without the proposed denoising approaches.
Our tokenizer implementation.
We implement our tokenizer using ViTs for both encoder and decoder [61, 18]. We adopt recent architectural advances from LLaMA [60], including RoPE [57] (together with learned positional embeddings following [22]), RMSNorm [71], and SwiGLU-FFN [53]. The encoder operates at a patch size of $16$, yielding $256$ latent tokens for each $256\times 256$ image, while the decoder uses patch size $1$ as there is no resolution change. The latent dimension is set to $16$. We omit the [CLS] token for simplicity.
Tokenizer training. Our tokenizer is trained using the weighted loss defined in Eq. 6, with default weights $\lambda_{\text{KL}}=10^{-6}$, $\lambda_{\text{percep}}=1.0$, and $\lambda_{\text{GAN}}=0.1$. In ablation studies, we use ViT-S for the encoder and ViT-B for the decoder, disable the GAN loss, and train for 50 epochs. We observe that including a GAN loss [66, 25, 19] sharpens reconstructions but roughly doubles training time, without altering result trends. For final experiments, we use ViT-B for both encoder and decoder, train for 200 epochs, and activate the GAN loss starting from epoch 100. All tokenizers use AdamW [41] with a global batch size of 1024 and peak learning rate of $4.0\times 10^{-4}$ (corresponding to a base learning rate of $1.0\times 10^{-4}$ scaled linearly by the global batch size divided by 256 [26]), linear warm-up, and cosine learning schedule. Further implementation details are provided in the Appendix.
Generative models.
To evaluate the broad effectiveness of a tokenizer, we experiment with six representative generative models, including three non-autoregressive models: DiT [46], SiT [42], and LightningDiT [64]; and three autoregressive models: MAR [36], causal RandomAR, and RasterAR based on RAR [68] and diffloss [36]. We follow the officially released implementations to reimplement all methods within a unified codebase, standardizing training and evaluation across methods. Previous works on visual tokenizers often exclusively evaluate on non-AR models (e.g., DiT, SiT), yet we find improvements from non-AR models do not necessarily translate to AR models (more on this later). Our experiments aim to provide useful data points toward a more universal tokenizer.
Generative model training.
We use a standardized training recipe for all generative models. Specifically, we follow the hyperparameters from [64], training generative models with a global batch size of 1024, using AdamW [41] with a constant learning rate of $2\times 10^{-4}$, without warm-up, gradient clipping, or weight decay. Autoregressive models adopt the three-layer 1024-channel diffloss MLP from [36]. For ablation studies, we train generative models for 100 epochs. For larger-scale experiments on MAR models, we train them for 800 epochs. All models utilize exponential moving average (EMA) with a decay rate of 0.9999. We always standardize tokenizer outputs by subtracting the mean and dividing by the standard deviation, both computed from the ImageNet training set. For publicly available tokenizers, we use their official standardization steps. Unless noted otherwise, we report FID@50k scores with classifier-free guidance (CFG), with optimal CFG scales searched from FID@10k results. We abbreviate FID@50k as FID.
5 Experiments
We now empirically analyze our proposed tokenizer, beginning with a series of ablation studies on noising strategies, followed by comprehensive comparisons across various generative modeling frameworks. Lastly, we benchmark against leading systems.
5.1 Main Properties
We study the generalizability of a tokenizer across non-autoregressive (non-AR) and autoregressive (AR) generative paradigms. To this end, we use SiT-B [42] and MAR-B [36] as representative non-AR and AR models, respectively. For all ablation studies, we adopt small encoders and base-sized decoders in our tokenizers.
5.1.1 Properties of Latent Noising
We first study latent noising as the sole form of deconstruction, without any masking.
Figure 2: Ablation on latent noise design. (a) Additive vs. interpolative noise. Interpolative noise clearly outperforms additive noise for both MAR [36] and SiT [42]. Both interpolative and additive latent noise lead to improved performance for MAR. (b) Latent noise standard deviation ($\gamma$). Our $l$-DeTok remains robust across various noise standard deviations. Generally, increasing $\gamma$ improves generation quality, with best results achieved around $\gamma=3.0$.
Interpolative vs. additive noise.
An important design choice in our $l$-DeTok is the use of interpolative latent noise instead of additive noise, which we ablate here. Specifically, we compare two latent noising variants: interpolative noise, i.e., $\mathbf{x}^{\prime}=(1-\tau)\mathbf{x}+\tau\boldsymbol{\varepsilon}$ (Eq. 4), and additive noise, i.e., $\mathbf{x}^{\prime}=\mathbf{x}+\tau\boldsymbol{\varepsilon}$. We set the noise standard deviation to $\gamma=1.0$ here. Figure 2-(a) presents the results.
Interpolative noise clearly outperforms additive noise for both SiT [42] and MAR [36]. This aligns with our expectation: interpolative noise explicitly ensures latent embeddings are heavily corrupted at high noise levels. In contrast, additive noise can potentially create shortcuts by making original signals remain dominant, reducing the effectiveness of the additive noise. Nonetheless, we observe the additive latent noise still improves generative performance for MAR but not for SiT. Understanding the underlying factors responsible for these model-specific differences would be an intriguing future direction, e.g., joint vs. factorized latent distribution estimation.
Noise standard deviation.
Figure 2-(b) studies the effect of noise standard deviation ($\gamma$ in Eq. 4). Both SiT and MAR consistently improve with interpolative latent noise across all tested levels. Performance peaks at moderately high standard deviation, indicating that stronger corruption generally yields more effective latents. This result confirms our key hypothesis: challenging denoising tasks naturally produce robust, downstream-aligned latents that benefit generative modeling.
5.1.2 Properties of Masking
We next investigate masking noise independently, without any latent noise.
Masking ratio.
We examine how varying the maximal masking ratio $M$ (Eq. 5) influences generation quality. As shown in Figure 3-(a), both SiT and MAR benefit from masking-based tokenizer training in generation quality. Masking ratios between 70% and 90% consistently yield stronger performance compared to low masking ratios (e.g., 30%), favoring high degrees of masking. This behavior mirrors observations from latent denoising, i.e., challenging denoising is more beneficial, indicating a common underlying principle under the deconstruction-reconstruction strategy.
Interestingly, SiT [42] still benefits from our masking-based $l$-DeTok despite not explicitly handling masked inputs during its own training. We hypothesize that masking implicitly promotes encoders to learn embeddings that are inherently robust other noise types, such as diffusion-based perturbations.
Constant vs. randomized masking ratio.
Figure 3-(b) compares randomized masking ratios against constant ones used in MAE [30]. For constant masking ratio experiments, we fine-tune the tokenizer decoder for an additional 10 epochs on full-visible latents to mitigate the distribution mismatch between training and inference, since fully visible inputs are absent during constant-ratio training. From Figure 3-(b), we see that randomized masking outperforms constant masking. Randomized masking encourages latent embeddings that are robust to varying corruption levels. This is more aligned with the downstream tasks, i.e., denoising across diverse levels of corruption.
[…]
5.1.3 Joint Denoising
Prior ablations indicate that both latent noising and masking can independently improve generation quality, with latent noising showing a stronger effect. Here, we investigate the effect of joint denoising. Based on prior results, we fix the noise standard deviation to $\gamma=3.0$ and the masking ratio to $M=0.7$. Figure 5 and Table 5 summarize the results.
Figure 4: Impact of tokenizer training strategies on generative performance. We compare our baseline tokenizer with our $l$-DeTok variants: masking-only ($M{=}0.7$), latent noising-only ($\gamma{=}3.0$), and their combination (joint noising). Both masking and latent noising independently improve generation quality, with latent noising showing a stronger effect. Joint noising further improves performance for MAR, particularly in inception scores (IS), but provides limited additional benefit for SiT when latent noising is already applied. FID@50k scores are detailed in Tab. 5.
Setup MAR-B SiT-B
w/ CFG w/o CFG w/ CFG w/o CFG
[…]
Baseline 3.31 247.63 16.70 6.97 181.61 26.82
Masking only 2.90 243.00 12.57 6.43 189.24 22.50
Latent noise only 2.77 249.02 8.82 5.56 193.52 16.11
Joint noise 2.65 263.03 7.48 5.50 195.07 16.03
+Extended 2.43 266.50 6.72 5.13 207.35 14.81
Figure 5: Effectiveness of denoising. We report FID and IS evaluated on 50,000 images here. Compared to baselines, we see substantial gains in generative models when using our $l$-DeTok. Extended: larger encoder, longer training, GAN enabled midway through training.
With joint denoising, our $l$-DeTok achieves FID scores of 5.50 (SiT-B) and 2.65 (MAR-B) with CFG (Table 5). In comparison, our baseline tokenizer—trained with identical settings but without any noise—obtains significantly worse results: 6.97 (SiT-B) and 3.31 (MAR-B). We observe that joint denoising is more effective for MAR, further pushing both FID and IS, but provides limited additional benefit for SiT when latent noising is already applied. This indicates that latent denoising is essential, while the masking-based denoising can be optional.
Lastly, with joint denoising, we increase the encoder to base size, train for 200 epochs, and enable the GAN loss starting from epoch 100 (+Extended in Table 5). Under this setting, performance improves to 5.13 (SiT) and 2.43 (MAR). We adopt this improved tokenizer for all subsequent evaluations.
5.2 Generalization Experiments
To comprehensively evaluate tokenizer generalizability, we compare performance across six representative generative models—three non-autoregressive (DiT [46], SiT [42], LightningDiT [64]) and three autoregressive (MAR [36], RandomAR, RasterAR)—using publicly available tokenizers introduced before. RandomAR and RasterAR are Transformer-based decoder-only (causal attention) variants adapted from RAR [68]. We modify them to support decoding continuous tokens via diffloss MLPs [36]. RandomAR generates tokens in random order while RasterAR uses a raster-scan order. We use base-sized models and train them for 100 epochs for experiments here.
Comparisons with standard convolutional tokenizers.
[…]
Comparisons with semantics-distilled tokenizers.
Our $l$-DeTok generalizes significantly better than semantics-distilled tokenizers. Table 1 also compares our method against recent semantics-distilled tokenizers such as VA-VAE [64] and MAETok [6], which distill semantics features from powerful pretrained encoders [43, 49]. Surprisingly, we empirically find that these tokenizers, despite their promising performance for non-AR models, do not generalize well to AR models. Specifically, FID scores (with CFG) degrade dramatically from 3.71, 11.78, and 7.99 to 16.66, 38.13, and 15.88 for MAR, RandomAR, and RasterAR, respectively. Previous studies implicitly presume that tokenizer improvements observed in one generative paradigm naturally extend to others. Yet, our experiments challenge this assumption, revealing a previously unrecognized gap in tokenizer transferability: tokenizer effectiveness in one generative paradigm does not necessarily transfer to others.
In sharp contrast, our method generalizes significantly better across both non-AR and AR models. Importantly, our $l$-DeTok requires no semantics distillation from large-scale pretrained encoders yet considerably surpasses standard tokenizer competitors. This lifts external dependencies and highlights a new and previously overlooked direction in tokenizer research.
Table 1: Generalizability comparison of tokenizers across different generative models. We compare various tokenizers on representative generative models. Our $l$-DeTok tokenizer outperforms other tokenizers for AR models, and also surpasses standard tokenizers trained without semantics distillation for non-AR models. All results are obtained with optimal CFG scales.
[…]
MAR RandomAR RasterAR SiT DiT Light.DiT
[…]
Tokenizers trained with semantics distillation from external pretrained models
VA-VAE [64] 0.28 16.66 144.5 38.13 68.3 15.88 160.5 4.33 221.1 4.91 213.9 2.86 275.1
MAETok [6] 0.48 6.99 201.8 24.83 97.6 15.92 127.2 4.77 243.2 5.24 224.7 3.92 273.3
Tokenizers trained without semantics distillation
SD-VAE [51] 0.61 4.64 259.8 13.11 141.8 8.26 179.3 7.66 187.5 8.33 179.8 4.24 223.7
MAR-VAE [36] 0.53 3.71 265.3 11.78 147.9 7.99 189.7 6.26 177.5 8.20 171.8 3.98 218.7
Our $l$-DeTok 0.68 2.43 266.5 5.22 248.9 4.46 257.7 5.13 207.3 6.58 173.9 3.63 225.4
5.3 Benchmarking with Previous Systems
We compare against leading generative systems in Table 2. For this experiment, we train MAR-B and MAR-L for 800 epochs. Simply adopting our tokenizer, without altering the MAR architecture, substantially improves the generative performance: MAR-B achieves an FID of 1.55 (from 2.31), and MAR-L further improves to 1.35 (from 1.78). Notably, our MAR-B and MAR-L both match or surpass the previously best-performing huge-size MAR model (1.35 vs. 1.55). Qualitative results are provided in Figure 6.
Table 2: System-level comparison on ImageNet 256$\times$256 class-conditioned generation. Our approach enables MAR models [36] to achieve leading results without relying on semantics distillation. : With additional decoder fine-tuning (see Sec. A.1 for details).
w/o CFG w/ CFG
[…]
With semantics distillation from external pretrained models
SiT-XL + REPA [70] 675M 5.90 157.8 0.70 0.69 1.42 305.7 0.80 0.64
SiT-XL + MAETok [6] 675M 2.31 216.5 - - 1.67 311.2 - -
LightningDiT + MAETok [6] 675M 2.21 208.3 - - 1.73 308.4 - -
LightningDiT + VAVAE [64] 675M 2.17 205.6 0.77 0.65 1.35 295.3 0.79 0.65
DDT-XL [63] 675M 6.27 154.7 0.68 0.69 1.26 310.6 0.79 0.65
Without semantics distillation from external pretrained models
ADM [16] 554M 10.94 101.0 0.69 0.63 4.59 186.7 0.82 0.52
VDM+++++ + [32] 2B 2.40 225.3 - - 2.12 267.7 - -
MaskGIT [5] 227M 6.18 182.1 0.80 0.51 - - - -
MAGVIT-v2 [67] 307M 3.65 200.5 - - 1.78 319.4 - -
LDM-4 [51] 400M 10.56 103.5 0.71 0.62 3.60 247.7 0.87 0.48
DiT-XL/2 [45] 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
SiT-XL/2 [42] 675M 8.30 - - - 2.06 270.3 0.82 0.59
MDTv2-XL/2 [23] 676M 5.06 155.6 0.72 0.66 1.58 314.7 0.79 0.65
MaskDiT [23] 675M 5.69 178.0 0.74 0.60 2.28 276.6 0.89 0.61
VAR-d30 [59] 2.0B - - - - 1.92 323.1 0.82 0.59
LlamaGen-3B  [58] 3.1B - - - - 2.18 263.3 0.81 0.58
RandAR-XXL  [44] 1.4B - - - - 2.15 322.0 0.79 0.62
FlowAR-H [50] 1.9B - - - - 1.65 296.5 0.83 0.60
CausalFusion [14] 676M 3.61 180.9 0.75 0.66 1.77 282.3 0.82 0.61
MAR-B + MAR-VAE [36] 208M 3.48 192.4 0.78 0.58 2.31 281.7 0.82 0.57
MAR-L + MAR-VAE [36] 479M 2.60 221.4 0.79 0.60 1.78 296.0 0.81 0.60
MAR-H + MAR-VAE [36] 943M 2.35 227.8 0.79 0.62 1.55 303.7 0.81 0.62
MAR-B [36] + our $l$-DeTok 208M 2.79 195.9 0.80 0.60 1.61 289.7 0.81 0.62
MAR-B [36] + our l-DeTok 208M 2.94 195.5 0.80 0.59 1.55 291.0 0.81 0.62
MAR-L [36] + our $l$-DeTok 479M 1.84 238.4 0.82 0.60 1.43 303.5 0.82 0.61
MAR-L [36] + our l-DeTok 479M 1.86 238.6 0.82 0.61 1.35 304.1 0.81 0.62
Figure 6: Qualitative Results. We show selected examples of class-conditional generation on ImageNet 256$\times$256 using MAR-L [36] trained with our tokenizer.
6 Discussion and Conclusion
Limitations. On the general side, our study primarily investigates continuous-valued tokenizers; the effectiveness of our denoising strategy for discrete, vector-quantized (VQ) tokenizers remains an open question. Exploring denoising-based objectives in the context of VQ tokenizers is a valuable future direction. Additionally, our evaluation is currently limited to the ImageNet benchmark. Assessing tokenizer scalability and generalization on broader, real-world datasets is an important next step. On the technical side, we observe a training/inference discrepancy in our tokenizer: the decoder is primarily trained on noise-injected latent embeddings, whereas it operates on almost noise-free embeddings during inference. Fine-tuning the decoder on clean latent embeddings partially addresses this discrepancy. Further investigation into mitigating this discrepancy could yield additional improvements. Nonetheless, the core motivation and insights of our work remain robust.
Conclusion. The strong generalization and effectiveness of our $l$-DeTok across different generative models highlights a fundamental yet underexplored opportunity: tokenizer representations alone can substantially advance different generative approaches without architectural changes. Our findings suggest that explicitly aligning tokenizer training with downstream denoising tasks is surprisingly beneficial for generative modeling, complementing traditional focuses on pixel-level accuracy or semantic alignment. Given the foundational role of tokenizers, we hope this perspective will inspire further research into effective and generalizable representation learning for generative models. Lastly, the core idea of this work is general and extends beyond image tokenization. Exploring its potential in broader generative modeling, such as video generation, action generation, protein design, and other domains, will be an exciting direction for future research.
Acknowledgment
The USC Geometry, Vision, and Learning Lab acknowledges generous supports from Toyota Research Institute, Dolby, Google DeepMind, Capital One, Nvidia, and Qualcomm. Yue Wang is also supported by a Powell Research Award.
Appendix
Appendix A Training and Inference Details
A.1 Tokenizer
Model.
Our tokenizers are based on ViTs [18]. We provide detailed model parameters in Table A.1. We include the details provided in Section 4 here for completeness: We implement our tokenizer using ViTs for both encoder and decoder [61, 18]. We adopt recent architectural advances from LLaMA [60], including RoPE [57] (together with learned positional embeddings following [22]), RMSNorm [71], and SwiGLU-FFN [53]. The encoder operates at a patch size of $16$, yielding $256$ latent tokens for each $256\times 256$ image, while the decoder uses patch size $1$ as there is no resolution change. The latent dimension is set to $16$. We omit the [CLS] token for simplicity.
Table A.1: Model Size. We report the model configurations and parameter counts of tokenizer encoders and decoders. The total tokenizer size is the sum of encoder and decoder parameters. For example, a tokenizer combining a ViT-S encoder and ViT-B decoder (S-B) has 111.6M parameters, while a tokenizer with both ViT-B encoder and decoder (B-B) has 171.7M parameters.
Size Hidden Size Blocks Heads Parameters
Small (S) 512 8 8 25.75M
Base (B) 768 12 12 85.85M
Training.
For ablation studies, we use a ViT-S encoder and a ViT-B decoder, disable GAN loss, and train for 50 epochs. For our final experiments, we adopt ViT-B for both encoder and decoder, enable GAN loss from epoch 100, and train for 200 epochs. In both settings, the global batch size is 1024, and the peak learning rate is set to $4.0\times 10^{-4}$ (scaled linearly from $1.0\times 10^{-4}$ at a batch size of 256). We apply linear warm-up for 25% of the total epochs (12 epochs for 50-epoch training, and 50 epochs for 200-epoch training), followed by cosine learning rate decay. We use the AdamW optimizer with $\beta$ parameters (0.9, 0.95) and a weight decay of $1.0\times 10^{-4}$. The only data augmentation employed is horizontal flipping. Our reconstruction loss closely follows the implementation in [69].
Training an S-B tokenizer (see Table A.1) without GAN loss for 50 epochs takes roughly 160 NVIDIA A100 GPU hours (enabling GAN loss from epoch 20 extends training to about 288 A100 GPU hours). Training a B-B tokenizer with GAN loss enabled from epoch 100 for a total of 200 epochs takes approximately 1,150 A100 GPU hours.
Decoder fine-tuning.
We observe an interesting discrepancy between training and inference in our $l$-DeTok. Our decoder is trained predominantly with noise-corrupted latent embeddings but encounters nearly clean embeddings during inference. To address this gap, we fine-tune the decoder on clean latent embeddings, i.e., masking and latent noising are disabled, for an additional 100 epochs. This adjustment partially alleviates the discrepancy, improving the FID of the 800-epoch MAR-L model from 1.43 to 1.35 and MAR-B model from 1.61 to 1.55 (Table 2). Figure A.2 and Figure A.1 compare the denoising capability of different tokenizers. The fine-tuned decoder exhibits reduced denoising ability.
Figure A.1: Visualization of latent denoising. Images generated from latent embeddings corrupted with varying noise levels ($t$) by the original decoder (top), fine-tuned decoder (middle), and baseline decoder (bottom). The fine-tuned decoder shows a reduced ability to recover images from noisy embeddings compared to the original decoder. The baseline tokenizer trained without the denoising objective fails to reconstruct original images from noisy latents.
Figure A.2: Visualization of “mask” denoising. Images generated from masked inputs with varying masking ratios ($m$) by the original decoder (top) and fine-tuned decoder (bottom). The fine-tuned decoder exhibits diminished capability in reconstructing masked regions compared to the original decoder.
Importantly, we find that the performance improvements from our approach primarily result from better latent representations rather than increased denoising capability of the decoder. To validate this, we perform an additional experiment by fine-tuning only the MAR-VAE decoder on the denoising task—keeping the encoder frozen, thus leaving latent representations unchanged. Contrary to expectations, this experiment leads to degraded performance, indicating that merely enhancing the decoder’s denoising capability exacerbates the training-inference discrepancy, since the generative models already excel at this task. These observations confirm that the key advantage of our method lies in improved latent representations, rather than decoder-side denoising improvements.
Pseudo-code of latent denoising
. See Algorithm 1.
Algorithm 1 Latent Denoising: PyTorch-like Pseudo-code
def denoise(x, encoder, decoder, max_mask_ratio=0.7, gamma=3.0):
# encode input image to latent embeddings under (optional) masking
z, ids_restore = encoder(x, max_mask_ratio=max_mask_ratio)
# variational latent embeddings
posteriors = diagonal_gaussian_dist(z)
z_sampled = posteriors.sample()
# sample interpolation factor uniformly from [0, 1]
bsz, n_tokens, chans = z_sampled.shape
device = z_sampled.device
noise_level = torch.rand(bsz, 1, 1, device=device).expand(-1, n_tokens, chans)
# generate Gaussian noise
noise = gamma * torch.randn(bsz, n_tokens, chans, device=device)
# interpolate latent embeddings with noise
z_noised = (1 - noise_level) * z_quantized + noise_level * noise
# reconstruct the inputs
recon = decoder(z_noised, ids_restore)
return recon
A.2 Generative Models
Model.
We re-implement all generative models within our codebase to standardize training and evaluation. We introduce one modification: for DiT [46], SiT [42], and LightningDiT [64], we apply classifier-free guidance (CFG) to all latent channels, rather than only the first three channels used in their original implementations. [^1] See original implementations in DiT, SiT, and LightningDiT. For training on 1D tokens (e.g., MAETok [6]), we use simple 1D learnable positional embeddings and disable RoPE used in LightningDiT. We adopt the default samplers from the original implementations, using 250 denoising steps during sampling.
For autoregressive (AR) methods, we follow the default MAR model and implement RandomAR and RasterAR following RAR [68]. To produce continuous tokens, we employ a diffloss [36] with a 3-layer, 1024-channel MLP head. We disable dropout in all MLP layers within Transformer blocks for autoregressive models. For inference, we use 64 autoregressive steps and 100 denoising steps for all experiments, except those in system-level comparison (Table 2), where we use 256 autoregressive steps. Sampling is performed with the default MAR sampler across all AR models. Following MAR [36], we set the sampling temperature to 1.0 when using CFG, and sweep temperatures when CFG is disabled (i.e., CFG=1.0).
Training.
We follow the training recipe from [64] for all experiments reported in this paper: global batch size of 1024, AdamW optimizer, constant learning rate of $2\times 10^{-4}$, and no warm-up, gradient clipping, or weight decay. We do not tune these hyperparameters.
Training DiT-B, SiT-B, and LightningDiT-B for 100 epochs takes approximately 128 to 200 A100 hours using locally cached tokens. Training MAR-B, RandomAR-B, and RasterAR-B for 100 epochs takes approximately 220 to 250 A100 hours. For the final MAR-B model used in the system-level comparison (Table 2), training for 800 epochs takes roughly 2,450 A100 hours; training MAR-L for the same duration takes about 3,850 A100 hours (online evaluation time included).
Appendix B Additional Results
Figure B.1: Tokenizer comparison for autoregressive models. We report FID@10k and IS@10k scores at different classifier-free guidance (CFG) scales. Semantics-distilled tokenizers (denoted by “*”) are shown in warm colors, while those without distillation are shown in cool colors. Our $l$-DeTok consistently achieves superior FID and IS metrics compared to other tokenizers. See Table 1 for results on 50,000 images evaluated at the optimal CFG scales.
FID vs. CFG curves.
Figures B.1 and B.2 compare how classifier-free guidance (CFG) scales influence generative performance (FID and IS) across different tokenizers and generative models. We use warm colors to denote semantics-distilled tokenizers (marked by “*”), and cool colors for tokenizers without distillation. Our $l$-DeTok consistently achieves stronger performance across varying CFG scales, improving notably over standard tokenizers and matching or even surpassing semantics-distilled ones. Final FID scores computed over 50,000 generated images at optimal CFG scales are summarized in Table 1.
Figure B.2: Tokenizer comparison for non-autoregressive models. We report FID@10k and IS@10k scores at different classifier-free guidance (CFG) scales. Semantics-distilled tokenizers (denoted by “*”) are shown in warm colors, while those without distillation are shown in cool colors. Our $l$-DeTok consistently outperforms standard (non-semantics-distilled) tokenizers, matching the performance of the best semantics-distilled tokenizer (VA-VAE [64]) and surpassing MAETok [6] in IS. See Table 1 for results on 50,000 images evaluated at the optimal CFG scales.